# EAGLE: Efficient Adaptive Geometry-based Learning in Cross-view Understanding

### v1

@@ 2 Related Work
Unsupervised Domain Adaptation Adversarial learning [6, 54, 55] and self-supervised learning [1, 65, 23, 14] are common approaches to UDA in semantic segmentation. The adversarial learning approaches are typically simultaneously trained on source and target data [54, 7, 6]. Chen et al. [7] first introduced an adversarial framework to domain adaptation. Later, several approaches improved adversarial learning by utilizing generative models [70, 34, 21], using additional labels [28, 55], incorporating with entropy minimization [54, 61, 50], or adopting the curriculum training [36]. Recently, the self-supervised approaches [1, 65, 23, 14] have achieved outstanding performance. Araslanov et al. [1] first proposed a self-supervised augmentation consistency framework for UDA. Hoyer et al. [23] utilized Transformers to improve the UDA performance. Later, this approach was further improved by utilizing multi-resolution cropped images [24] and masked image consistency strategy [25] to enhance contextual learning. Recent studies improved the self-supervised approach by aligning both output and attention levels via the cross-domain prediction consistency framework [56], using a prototypical representation [65], learning the cross-model consistency via depths [63], improving the class-relevant fairness [51], or exploring the relations of pseudo-labels [67]. Fashes et al. [15] introduced a prompt-based feature augmentation method to zero-shot UDA. Gong et al. [18] introduced a geodesic flow kernel to model the manifold structure between domains. Later, Simon et al. [46] designed distillation loss by the geodesic flow path.
@@
Vision-Language and Open-Vocab Segmentation By pre-training on a large-scale vision-language dataset [38, 27], the vision-language models can learn various visual concepts and can further be transferred to other vision problems through “prompting[17, 37, 31], e.g., open-vocab segmentation [60, 13, 37]. Li et al. [29] first introduced the language-driving approach to semantic segmentation. Rao et al. [39] represented a context-aware prompting mechanism for dense prediction tasks. Ghiasi et al. [17] proposed an OpenSeg framework that learns the visual-semantic alignments. Qin et al. [37] presented a unified, universal, and open-vocab segmentation network based on Mask2Former [8] with an adaptive prompting mechanism. Xu et al. [60] proposed a two-stage open-vocab segmentation framework using the mask proposal generator and the pre-trained CLIP model. Ding et al. [13] decoupled the zero-shot semantic segmentation to class-agnostic segmentation and segment-level zero-shot classification. Liang et al. [31] improved the two-stage open-vocab segmentation model by further fine-tuning CLIP on masked image regions and corresponding descriptions.
@@
Cross-view Learning The early studies exploited cross-view learning in geo-localization by using a polar transform across views [45, 44] or generative networks to cross-view images [40, 48]. Meanwhile, Zhu et al. [71] exploited the correlation between street- and aerial-view data via self-attention. In semantic segmentation, Coors et al. [10] first introduced a cross-view adaptation approach utilizing the depth labels and the cross-view transformation between car and truck views. However, this change of views in [10] is not as big a hurdle as the change of views in our problem, i.e., car view to drone view. Ren et al. [41] presented an adaptation approach across viewpoints using the 3D models of scenes to create pairs of cross-view images. Vidit et al. [53] modeled the geometric shift in cross FoV setting for object detection by learning position-invariant homography transform. Di Mauro et al. [12] introduced an adversarial method trained on a multi-view synthetic dataset where images are captured from different pitch and yaw angles at the same altitudes of the camera positions. Meanwhile, in our problem, the camera views could be placed at different altitudes (e.g., the car and the drone), which reveals large structural differences between the images. Truong et al. [49, 52] first introduced a simple approach to model the relation across views. However, these methods lack a theory and a mechanism for cross-view geometric structural change modeling. To the best of our knowledge, there are limited studies exploiting cross-view adaptation in semantic segmentation. Therefore, our work presents a new approach to model the geometric correlation across views.
@@ 4.2 Ablation Study
4.3 Comparisons with Prior UDA Methods
SYNTHIA $\to$ UAVID As shown in Table 4, our EAGLE has achieved SOTA results and outperforms prior UDA methods by a large margin. For fair comparisons, we adopt the DeepLab [3] and DAFormer [23] for the segmentation network. In particular, our mIoU results using DeepLab and DAFormer are $45.2\%$ and $50.8\%$. In the DAFormer backbone, the mIoU results of our approach are higher than CROVIA [49] and MIC [25] by $+4.8\%$ and $+9.0\%$. The IoU result of each class also consistently outperformed the prior methods. Highlighted that although our approach does NOT use depth labels, our results still outperform the one using depths, i.e., DADA [55]. It has emphasized that our approach is able to better capture the cross-view structural changes compared to prior methods.
GTA $\to$ UAVID As shown in Table 4, our effectiveness outperforms prior domain adaptation approaches when measured by both mIoU performance and the IoU accuracy of each class. In particular, our mIoU performance using DeepLab and DAFormer network achieves $36.7\%$ and $40.7\%$, respectively. Our results have substantially closed the performance gap with the supervised results. By using the better segmentation-based network, i.e., Mask2Former with ResNet, the performance of our approach is further improved to $40.1\%$ compared to DeepLab with ResNet.
@@ 4.4 Comparisons with Open-vocab Segmentation
Open-vocab Segmentation on Unseen Classes Table 6 illustrates the experimental results of our cross-view adaptation approach on unseen classes. In this experiment, we consider classes of Tree and Person as the unseen classes. As shown in the results, our cross-view adaptation approach with a view-condition prompting mechanism has achieved the best mIoU performance on unseen classes on both benchmarks, i.e., $39.3\%$ and $39.6\%$ on two benchmarks. Our experimental results have further confirmed the effectiveness and the generalizability of our cross-view geometric modeling and view-condition prompting approach to the open-vocab segmentation across camera views.
Real-to-Real Cross-view Adaptation Setting We evaluated our approach in the real-to-real setting, i.e., BDD $\to$ UAVID. Our approach is evaluated in two different settings, i.e., Unsupervised Domain Adaptation and Open-Vocab Semantic Segmentation. As shown in Table 7, our results have shown a significant improvement in our approach in real-to-real settings in both unsupervised domain adaptation and open-vocab semantic semantic segmentation. While the results of prior unsupervised domain adaptation, i.e., BiMaL [50] and DAFormer [23], gain limited performance due to their limits in cross-view learning, our method outperforms other methods these prior methods by a large margin.
@@ 3 Ablation Study
Subspace Representation of Geodesic Flow-based Metrics To illustrate the ability of structural learning of our geodesic flow-based metrics, we use a subset of images of the car-view and the drone-view dataset to visualize the base structure of subspaces obtained from the PCA algorithm. Fig. 5 visualizes the mean structures of car-view and drone-view images. As shown in Fig. 5, The subspaces of car-view images represent the geometric structures of car-view data, i.e., the road in the middle, buildings, trees on two sides, etc. Meanwhile, the geometric structures of the drone view have also been illustrated in the figure with structures and topological distributions of objects (e.g., the road in the middle and trees and buildings on the sides) on the scenes. The results have illustrated the base geometric structures of the car-view and the drone-view data. Then, by modeling the geodesic flow path across two subspaces, our metric is able to measure the cross-view geometric structural changes (i.e., the change of structures and topological layouts of the scene) from the car view to the drone view. Our experimental results in other ablation studies have further confirmed our effectiveness in geometric structural modeling across views.

### v2

# EAGLE [Uncaptioned image]: Efficient Adaptive Geometry-based Learning in Cross-view Understanding

@@ 2 Related Work
Unsupervised Domain Adaptation Adversarial learning [6, 58, 59] and self-supervised learning [1, 69, 23, 14] are common approaches to UDA in semantic segmentation. The adversarial learning approaches are typically simultaneously trained on source and target data [58, 7, 6]. Chen et al. [7] first introduced an adversarial framework to domain adaptation. Later, several approaches improved adversarial learning by utilizing generative models [74, 34, 21], using additional labels [28, 59], incorporating with entropy minimization [58, 65, 51, 52], or adopting the curriculum training [37]. Recently, the self-supervised approaches [1, 69, 23, 14] have achieved outstanding performance. Araslanov et al. [1] first proposed a self-supervised augmentation consistency framework for UDA. Hoyer et al. [23] utilized Transformers to improve the UDA performance. Later, this approach was further improved by utilizing multi-resolution cropped images [24] and masked image consistency strategy [25] to enhance contextual learning. Recent studies improved the self-supervised approach by aligning both output and attention levels via the cross-domain prediction consistency framework [60], using a prototypical representation [69], learning the cross-model consistency via depths [67], improving the class-relevant fairness [53, 55, 56], or exploring the relations of pseudo-labels [71]. Fashes et al. [15] introduced a prompt-based feature augmentation method to zero-shot UDA. Gong et al. [18] introduced a geodesic flow kernel to model the manifold structure between domains. Later, Simon et al. [47] designed distillation loss by the geodesic flow path.
@@
Vision-Language and Open-Vocab Segmentation By pre-training on a large-scale vision-language dataset [39, 27], the vision-language models can learn various visual concepts and can further be transferred to other vision problems through “prompting[17, 38, 31, 35], e.g., open-vocab segmentation [64, 13, 38]. Li et al. [29] first introduced the language-driving approach to semantic segmentation. Rao et al. [40] represented a context-aware prompting mechanism for dense prediction tasks. Ghiasi et al. [17] proposed an OpenSeg framework that learns the visual-semantic alignments. Qin et al. [38] presented a unified, universal, and open-vocab segmentation network based on Mask2Former [8] with an adaptive prompting mechanism. Xu et al. [64] proposed a two-stage open-vocab segmentation framework using the mask proposal generator and the pre-trained CLIP model. Ding et al. [13] decoupled the zero-shot semantic segmentation to class-agnostic segmentation and segment-level zero-shot classification. Liang et al. [31] improved the two-stage open-vocab segmentation model by further fine-tuning CLIP on masked image regions and corresponding descriptions.
@@
Cross-view Learning The early studies exploited cross-view learning in geo-localization by using a polar transform across views [46, 45] or generative networks to cross-view images [41, 49]. Meanwhile, Zhu et al. [75] exploited the correlation between street- and aerial-view data via self-attention. In semantic segmentation, Coors et al. [10] first introduced a cross-view adaptation approach utilizing the depth labels and the cross-view transformation between car and truck views. However, this change of views in [10] is not as big a hurdle as the change of views in our problem, i.e., car view to drone view. Ren et al. [42] presented an adaptation approach across viewpoints using the 3D models of scenes to create pairs of cross-view images. Vidit et al. [57] modeled the geometric shift in cross FoV setting for object detection by learning position-invariant homography transform. Di Mauro et al. [12] introduced an adversarial method trained on a multi-view synthetic dataset where images are captured from different pitch and yaw angles at the same altitudes of the camera positions. Meanwhile, in our problem, the camera views could be placed at different altitudes (e.g., the car and the drone), which reveals large structural differences between the images. Truong et al. [50, 54] first introduced a simple approach to model the relation across views. CROVIA [50] measures the cross-view structural changes by measuring the distribution shift and only focuses on the cross-view adaptation setting in semantic segmentation. However, these methods [50, 54] lack a theory and a mechanism for cross-view geometric structural change modeling. To the best of our knowledge, there are limited studies exploiting cross-view adaptation in semantic segmentation. Therefore, our work presents a new approach to model the geometric correlation across views.
@@ 4.2 Ablation Study
Figure 4: The Qualitative Results of Cross-View Adaptation (Without Prompt).
Figure 5: The Qualitative Results of Cross-View Adaptation (With Prompt).
Qualitative Results. To further illustrate the effectiveness of our proposed, we visualize the results produced by our model. In the model without prompting, Figure 4 illustrates the results of our cross-view adaptation compared to those without adaptation. As shown in the results, our approach can effectively segment the objects in the drone view. We also compare with the prior ProDA [69] and CROVIA [50] methods. Our qualitative results remain better than the prior adaptation method. For the model with prompting, Figure 5 illustrates the effectiveness of our approach in three cases: without adaptation, with cross-view adaptation, and with view-condition prompting. As shown in the results, our cross-view adaptation can efficiently model the segmentation of the view. By using the view-condition prompting, our model can further improve the segmentation of persons and vehicles.
4.3 Comparisons with Prior UDA Methods
@@
SYNTHIA $\to$ UAVID As shown in Table 4, our EAGLE has achieved SOTA results and outperforms prior view transformation (i.e., Polar Transform [45]) UDA methods by a large margin. For fair comparisons, we adopt the DeepLab [3] and DAFormer [23] for the segmentation network. In particular, our mIoU results using DeepLab and DAFormer are $45.2\%$ and $50.8\%$. In the DAFormer backbone, the mIoU results of our approach are higher than CROVIA [50] and MIC [25] by $+4.8\%$ and $+9.0\%$. The IoU result of each class also consistently outperformed the prior methods. Highlighted that although our approach does NOT use depth labels, our results still outperform the one using depths, i.e., DADA [59]. It has emphasized that our approach is able to better capture the cross-view structural changes compared to prior methods. Figure 4 illustrates our qualitative results compared to ProDA [69] and CROVIA [50].
GTA $\to$ UAVID As shown in Table 4, our effectiveness outperforms prior polar view transformation [45] and domain adaptation approaches when measured by both mIoU performance and the IoU accuracy of each class. In particular, our mIoU performance using DeepLab and DAFormer network achieves $36.7\%$ and $40.7\%$, respectively. Our results have substantially closed the performance gap with the supervised results. By using the better segmentation-based network, i.e., Mask2Former with ResNet, the performance of our approach is further improved to $40.1\%$ compared to DeepLab.
@@ 4.4 Comparisons with Open-vocab Segmentation
Open-vocab Segmentation on Unseen Classes Table 6 illustrates the experimental results of our cross-view adaptation approach on unseen classes. In this experiment, we consider classes of Tree and Person as the unseen classes. As shown in the results, our cross-view adaptation approach with a view-condition prompting mechanism has achieved the best mIoU performance on unseen classes on both benchmarks, i.e., $39.3\%$ and $39.6\%$ on two benchmarks. Our experimental results have further confirmed the effectiveness and the generalizability of our cross-view geometric modeling and view-condition prompting approach to the open-vocab segmentation across views.
@@
Real-to-Real Cross-view Adaptation Setting We evaluated our approach in the real-to-real setting, i.e., BDD $\to$ UAVID. Our approach is evaluated in two different settings, i.e., Unsupervised Domain Adaptation and Open-Vocab Semantic Segmentation. As shown in Table 7, our results have shown a significant improvement in our approach in real-to-real settings in both unsupervised domain adaptation and open-vocab semantic segmentation. While the results of prior unsupervised domain adaptation, i.e., BiMaL [51] and DAFormer [23], gain limited performance due to their limits in cross-view learning, our method outperforms other methods these prior methods by a large margin.
@@ 5 Conclusions
Acknowledgment This work is partly supported by NSF Data Science, Data Analytics that are Robust and Trusted (DART), NSF SBIR Phase 2, and Arkansas Biosciences Institute (ABI) grants. We also acknowledge the Arkansas High-Performance Computing Center for providing GPUs.
@@ 3 Ablation Study
Subspace Representation of Geodesic Flow-based Metrics To illustrate the ability of structural learning of our geodesic flow-based metrics, we use a subset of images of the car-view and the drone-view dataset to visualize the base structure of subspaces obtained from the PCA algorithm. Fig. 7 visualizes the mean structures of car-view and drone-view images. As shown in Fig. 7, The subspaces of car-view images represent the geometric structures of car-view data, i.e., the road in the middle, buildings, trees on two sides, etc. Meanwhile, the geometric structures of the drone view have also been illustrated in the figure with structures and topological distributions of objects (e.g., the road in the middle and trees and buildings on the sides) on the scenes. The results have illustrated the base geometric structures of the car-view and the drone-view data. Then, by modeling the geodesic flow path across two subspaces, our metric is able to measure the cross-view geometric structural changes (i.e., the change of structures and topological layouts of the scene) from the car view to the drone view. Our experimental results in other ablation studies have further confirmed our effectiveness in geometric structural modeling across views. Figure 8 illustrates the feature distributions with and without our proposed approach. As shown in Figure 8, our approach can help to improve the feature representations of classes, and the cluster of each class is more compact, especially in classes of car, tree, and person.
@@
Figure 8: The Feature Distribution of Classes in SYNTHIA $\to$ UAVID Experiments.